if 50.0ml of Ca(OH)2 is neutralized when titraties with 15.5ml of 0.0850 M HCl, what is the concentration of the Ca(OH)2?
1
Expert's answer
2020-08-06T03:57:53-0400

Ca(OH)2 + 2HCl = CaCl2 + 2H2O

n(HCl)= 0,085*0,0155=0,0013 mol

C(Ca(OH)2)= (0,0013/2)/0,05=0,013M


Answer: C (Ca(OH)2)=0,013M